Calculate the pH of the following buffer solutions:

(a) a solution made by dissolving 4.550 g KH2PO4 and 1.871 g K2HPO4 in water to give 350. mL of solution.

(b) a solution made by mixing 35.00 mL 0.35 M N2H4 with 15.00 mL 0.21 M N2H5Cl.

The base dissociation constant Kb is 8.50 x 10-7 for N2H4 . Assume that the final volume is the sum of the volunes of the two solutions that are mixed.
Answer #1

First calculate molarity of both KH2PO4 and K2HPO4 and then molarity equation M1V1+M2V2=M(V1+V2) and calculate pH.From formula M(V1+V2)=M1V1+M2V2, calculate M, then pH.